# One Queue Is All You Need: Resolving Head-of-Line Blocking in Large Language Model Serving

### v1

@@ 3.1 Lifecycle of a Request in QLM
To explain the design of QLM (as shown in Figure 6), we first walk through the lifecycle of a request generated from applications to completion. LLM-augmented applications generate requests that are received at the QLM API gateway. These requests are added to a global queue where they wait until being served. To prevent the global queue from being a single point of failure, it is implemented with a distributed message broker such as RabbitMQ [40] that provides the requisite fault tolerance and consistency properties.
@@
Formation of Request Groups. Each request is grouped with other requests that share common performance characteristics (such as model type, SLO value, and token distribution) to form Request Groups. This converts the complexity of the optimization problem from per-request level to per-request-group level. By doing so, it alleviates the scalability challenges and lowers optimization overheads. Request grouping criteria and details are described in section 4.
@@
Assigning Request Groups to Virtual Queues. Requests in a request group are then assigned to a Virtual Queue, representing a waiting queue for an LLM serving instance in the cluster. The introduction of virtual queues creates a common abstraction for setting the actions of backend LSOs. The ordering of the request groups in a virtual queue determines the execution ordering of the requests on the corresponding LLM serving instance. We refer readers to section 4 for virtual queue formation and section 5 for the translation from virtual queue ordering to LSO actions.
@@
Virtual Queue Reordering for SLO Attainment Maximization. While requests are assigned to virtual queues in a first-come-first-serve manner, request groups in a virtual queue are reordered to maximize the SLO attainment for all requests being served. At the core of SLO attainment maximization are QLM’s request completion time (RCT) estimator (see section 6) and plan generator (see section 7).
@@ 3.2 QLM Design Principles
Design Principle #1: Scaling to a high request arrival rate and burstiness. QLM must be able to handle a high volume of requests for SLO attainment without the overhead that compromises the serving throughput. Existing model-serving frameworks that leverage optimization techniques such as linear programming have exponential or cubical complexity, which limits the scalability to larger workloads and longer queues. QLM instead introduces request groups (see section 4) to reduce the input space of the optimization solvers, thus lowering the computational overhead and enabling scalability.
Design Principle #2: Handling a diverse set of LSOs with complex inter-relationships. In Section 2.4 (Insight #2 and Insight #3), we demonstrate that multiple LSOs can either conflict with each other or have a synergistic effect to enhance overall serving performance. To attain model-serving latency SLOs, it is thus critical to translate an end-to-end SLO to the appropriate backend LSO actions. QLM models these complex inter-relationships with a two-step approach. First, virtual queues (see section 4) enable the necessary abstraction to enable actions for multiple backend LSOs. Second, the plan generator (see section 7) models the impact of ordering on multiple LSOs with a stochastic programming solver. We specifically prefer a stochastic programming solver over other optimization methods because it: (1) allows us to represent the stochastic nature of LLM’s autoregressive generation, and (2) systematically considers various constraints introduced by multiple LSOs.
Design Principle #3: Handling heterogeneous models and hardware device configurations. LLM serving workloads consist of diverse model types with vastly different computational requirements, SLOs, and token length distributions. Hardware device configurations are also heterogeneous in terms of computing power, GPU memory capacity, and GPU-CPU memory bandwidth. To efficiently map LLM requests to the appropriate hardware resources, QLM’s plan generator has to consider each device’s computing power, memory capacity, and memory bandwidth. The RCT estimator (see section 6) estimates this impact of heterogeneity for the plan generator. The profiling costs for the RCT estimator are minimal, only a single batch run for a given combination of request group and GPU device is needed. Hence, QLM does not require significant training when adding new LLM models or GPU devices into the serving cluster.
4 Virtual Queues
In this section, we describe the concept of virtual queues and the process of classifying LLM requests into request groups and assigning request groups to virtual queues. QLM’s virtual queues draw inspiration from virtual output queuing (VOQ) [32, 47], a popular architecture used in network switches to address head-of-line blocking.
@@ 4 Virtual Queues
Request Group: Each request group is a collection of multiple requests that are relatively homogeneous, i.e., sharing similar performance demand or requirement characteristics. We identify that input/output token distributions, model type, and SLO values are sufficient for the RCT estimator (as explained in Section 6).
@@
By creating the abstraction of request groups and virtual queues, the ordering of request groups in a virtual queue allows QLM to configure actions for multiple downstream LSOs to attain end-to-end latency SLOs for the LLM-serving requests in a scalable manner (described in Section 5).
@@
Request Group Creation. Request groups are created in two steps: (i) clustering similar requests based on definition 4.1, and (ii) splitting large request groups. Algorithm 1 describes the request group creation process. The parameters identified for the request grouping include model types, input/output token distribution, and SLO values. Grouped requests based on such parameters exhibit predictable request completion time distribution (compared to that of each individual request) as explained in Section 6. Additionally, we also limit the size of each request group to a small multiple ($\delta$) of batch size. We refer the reader to section 8.3 for the trade-off analysis between overhead and decision-making granularity: (1) Larger request groups would decrease the number of request groups and thus the overhead of the plan generation; (2) However, restricting the size of request groups is beneficial as it allows for more fine-grained decisions. Since requests within a request group are relatively homogeneous, QLM treats the ordering of the requests within a group using a first-come-first-serve (FCFS) policy.
@@
Handling New Incoming Requests. As new requests join the global queue, they are classified into the existing request groups, and the RCT estimator calculation (section 6) is triggered to find out whether any SLOs are being violated. Upon any SLO violation, the plan generator (section 7) is called to reorder the request groups in the virtual queues to maximize SLO attainment given the current states (estimations).
@@
Fault Tolerance in Queue Management. QLM only stores a single replica of the requests and their metadata in the global queue, which avoids the need to maintain consistency between multiple queues. The global queue is implemented using a distributed message queue broker such as RabbitMQ [40] that provides the necessary replication, fault tolerance, and persistence mechanisms. The data structure that implements virtual queues records orderings of subsets of requests in the global queue. These virtual queues are implemented as lightweight data structures that maintain pointers or references to the actual requests stored in the global queue. By using virtual queues, QLM can achieve the following: (1) Fault Isolation: If an LLM serving instance fails, only the corresponding virtual queue is affected, and the remaining virtual queues can continue processing requests without interruption. Request groups from the lost virtual queue are assigned to other virtual queues using the plan generator. (2) Consistency: Since the actual requests are stored in the global queue, virtual queues can be reconstructed or reassigned without compromising the consistency of the request data.
5 LLM Serving Operations
@@
The LLM serving instances serve requests from the corresponding virtual queue and execute backend LSO actions when necessary. The LSOs by themselves are merely action actuators, and the intelligence required to configure when and which action to set comes from the virtual queue ordering set by the plan generator (as described in section 7).
@@
fig. 7 shows the five basic LSOs that QLM currently supports. A QLM agent resident on each LLM serving instance monitors the virtual queue ordering and converts it into LSO actions. When the virtual queue state changes or new requests are added, QLM agents initiate request pulling and load balancing. Similarly, when the head request group changes, the QLM agent initiates request eviction or GPU-CPU state swap. Model warm start is initiated by the QLM agent for models at the head of the virtual queue. Each of these LSOs modifies the internal state of the LLM serving instance, which includes the running batch of requests, KV cache store, and model weights. Below, we describe each of these LSOs in detail and their action setup based on the virtual queue ordering.
@@
Figure 7: Basic LLM serving operations (LSOs) for an LLM-serving instance that a QLM agent manages.
@@
Request Eviction ( 2 in fig. 7). As request pulling by itself may not be sufficient to enable the immediate serving of requests with low SLO values due to head-of-the-line blocking, QLM also supports request eviction. Request eviction is invoked when the RCT estimator detects an SLO violation, and the plan generator replaces an existing request group by placing a request group at the head of the virtual queue. In request eviction, requests of the head request group are pulled into the running batch based on available capacity, and previously running requests are evicted (back) into the global queue. As each request is evicted, it loses its KV cache, which needs to be regenerated in the future. However, the regeneration process is faster than the original prompt generation as the output tokens generated so far are appended to the input prompt to generate the remaining output tokens. Therefore, the original prefill and multiple decode steps are replaced by a single prefill step.
@@
GPU-CPU State Swapping ( 3 in fig. 7). Request eviction requires repeating the prefill step when re-executing the evicted requests, leading to repetitive and redundant computation. The prefill step could be costly for less powerful GPU hardware or complicated model types. An alternative to request eviction is to migrate the KV cache of evicted requests into the CPU memory. However, the GPU-to-CPU memory bandwidth is typically at least 10$\times$ less than the GPU memory bandwidth, and if the evicted request has a large KV cache, it leads to significant transfer overhead and consequent performance degradation. QLM hides this performance degradation using the asynchronous GPU memory copy available in most GPU programming libraries. As an alternative to request eviction, GPU-CPU state swapping can also be invoked when the head request group is updated.
@@
Load Balancing ( 4 in fig. 7). As each LLM serving instance is associated with a separate virtual queue, the plan generator’s assignment of request groups to a virtual queue inherently performs load balancing. Each instance would only pull from its associated virtual queue, thus ensuring the distribution of requests across all the serving instances. Note that QLM does not implement preemptive load balancing, i.e., once request groups start executing on an LLM serving instance, they cannot be migrated to another instance.
@@
Model Warm Start ( 5 in fig. 7). Each LLM serving instance can serve multiple models by switching the underlying model weights and flushing out the KV cache. QLM assumes a two-tier hierarchy of memory and disk storage. Therefore, any model that needs to be served from the LLM model registry (located in the storage) has to undergo two distinct swaps: (1) Storage-CPU swapping:The model is first swapped from the LLM model registry to CPU memory, and (2) CPU-GPU swapping:The model located in the CPU memory is swapped into GPU memory for inferencing. QLM is able to decide the state of each model by checking the virtual queue order. The model for the head request of the virtual queue is currently active and should be placed in the GPU memory. Models present later in the virtual queue are warm and placed in the CPU memory until all the CPU memory is exhausted. The remaining models (cold models) are not swapped out from the LLM model registry (located in the storage).
@@
LSO Implementation. We implement the abovementioned LSOs on top of vLLM, a state-of-the-art LLM serving system. QLM agents are responsible for triggering each LSO action. Request pulling and load balancing are implemented by async pull calls to the virtual queues when there is spare token capacity on the LLM serving instance. Request eviction and GPU-CPU state swapping require instrumentation to the vLLM scheduler. In each iteration, the vLLM scheduler attempts to generate a new token for all the running requests and preempts any request that exceeds the total GPU capacity. At the end of an iteration, QLM agent checks if either request eviction or state swapping is required and performs the operation. For both operations, current requests are removed from the running batch to make space for the incoming requests at the head of the virtual queue. For request eviction, the scheduler only saves the output tokens generated, and for state swapping, it uses an asynchronous GPU transfer of the KV cache. Model swapping into the GPU is implemented by changing the underlying model of the vLLM instance and flushing out the KV cache.
6 Request Completion Time (RCT) Estimator
Table 1: Glossary of symbols used in the RCT estimator.
@@ 6 Request Completion Time (RCT) Estimator
For requests within each request group, the RCT estimator leverages a Bayesian statistical approach to generate probability distributions of the completion times and waiting times. To create such RCT estimator, we extend the estimator described in Section 2.4 further by accounting for the variance per request in completion times. The process is explained below with variable definitions listed in table 1.
As shown in Equation 1, the total completion time equals the sum of the waiting time ($W_{q}$), prefill time ($P$), and total decode time across all the output tokens ($D_{q}$) for a request $q$.
$C_{q}=W_{q}+P+D_{q}$ (1)
Estimating Prefill Time. The prefill time $P$ is typically constant per model type as it is a highly parallel GPU-accelerated operation whose time increases minimally as the number of input tokens increases. Experiments show that the latency increase from additional input tokens is 100$\times$ less compared to the latency increase from each additional output token [2]. Therefore, the major distribution terms that still remain to be estimated are the waiting time ($W_{q}$) and decode time ($D_{q}$).
Estimating Waiting Time. As token generation throughput has a low variance due to continuous batching (characterized in Insight #1 Section 2.4), we consider the token generation throughput ($\Theta$) to be constant throughout the token generation process. Therefore, the total waiting time for a single request can be represented by Equation 2 by dividing the number of tokens ahead ($\sum_{i=1}^{q-1}O_{i}$) in the queue by the token generation throughput ($\Theta$) where $i$ denotes each of the $q-1$ requests in the queue ahead of the request we model.
$W_{q}=\sum_{i=1}^{q-1}\frac{O_{i}}{\Theta}$ (2)
Note that we do not know the number of output tokens ahead of time (that requires the knowledge of the output sequence for all requests in the waiting queue), so we model them as a normal distribution (as shown in Equation 3) with the mean $\mu_{o}$ and standard deviation $\sigma_{o}$ fitted from the offline dataset of the request input-output history for the request group that the request $q$ belongs to.
$O_{q}\sim N(\mu_{o},\sigma_{o})$ (3)
@@
$D_{q}=O_{q}\times\epsilon\times d$ (4)
If GPU memory was not a constraint, the decode steps would not be interrupted, and the total decode time would simply be the product of the number of output tokens ($O_{q}$) and decode time per output token ($d$). However, LLM serving systems cannot ensure this ideal behavior due to continuous batching. As requests are added continuously (due to iteration-level scheduling) to the GPU’s running batch, some requests inevitably exceed the total GPU memory capacity limit and have to be temporarily preempted. This leads to inefficiency in the generation process that we capture with the inefficiency factor $\epsilon$, i.e., a constant multiplied by the decode time per token that captures the inefficiency associated with the generation process.
@@
$C=\max_{q}C_{q}$ (5)
Offline Profiling. There are two independent profiling steps required for the RCT estimator: (a) Workload Profiling:samples multiple requests from the workload to generate a distribution for input and output tokens, and (b) Hardware Profiling:requires running the model with a single batch of requests on the specific GPU. Fixed variables associated with the model, such as the prefill time ($P$), decode time per iteration ($d$), and inefficiency factor ($\epsilon$), are obtained during hardware profiling by logging metrics from the LLM serving instance.
7 Plan Generator
The plan generator is invoked by the RCT estimator when an SLO violation is likely to occur. Upon invocation, the plan generator runs a stochastic programming model to reorder the virtual queues that decide underlying LSO actions to maximize SLO attainment. Details on how the queue ordering leads to the selection and actuation of LSO actions in each LLM serving instance are defined in Section 5. The plan generator uses a stochastic program solver because it: (a) allows handling non-determinism by representing request group completion times as distributions, and (b) offers a systematic way to model various constraints associated with SLOs, model swapping times, and hardware heterogeneity. In this section, we present the stochastic programming model with its defined variables listed in table 2.
Table 2: Glossary of symbols used in the stochastic programming solver.

### v2

@@ 3.1. Lifecycle of a Request in QLM
To explain the design of QLM (as shown in Figure 6), we first walk through the lifecycle of a request generated from applications to completion. LLM-augmented applications generate requests that are received at the QLM API gateway. These requests are added to a global queue where they wait until being served. To prevent the global queue from being a single point of failure, it is implemented with a distributed message broker such as RabbitMQ (RabbitMQ, 2024) that provides the requisite fault tolerance and consistency properties.
@@
Formation of Request Groups. Every incoming request is grouped with other requests that share common performance characteristics (such as model type, SLO value, and token distribution) to form Request Groups. This converts the complexity of the optimization problem from per-request level to per-request-group level. By doing so, it alleviates the scalability challenges and lowers optimization overheads. Additionally, request groups are a useful abstraction in the multi-model serving case as described in Section 2.4. Request grouping criteria and details are described in section 4.
@@
Assigning Request Groups to Virtual Queues. Requests in a request group are then assigned to a Virtual Queue, representing a waiting queue for an LLM serving instance in the cluster. The introduction of virtual queues creates a common abstraction for setting the actions of backend LLM Serving Operations (LSOs) such as request pulling, request eviction, load balancing, and model swapping. The ordering of the request groups in a virtual queue determines the execution ordering of the requests on the corresponding LLM serving instance. We refer readers to section 4 for virtual queue formation and section 5 for the translation from virtual queue ordering to LSO actions.
@@
Virtual Queue Reordering for SLO Attainment Maximization. While requests are assigned to request groups in a first-come-first-serve manner, request groups in a virtual queue are reordered to maximize the SLO attainment for all requests being served. At the core of SLO attainment maximization are QLM’s request waiting time (RWT) estimator (see section 6) and global scheduler (see section 7).
@@ 3.2. QLM Design Principles
Design Principle #1: Scaling to a high request arrival rate and burstiness. QLM must be able to handle a high volume of requests for SLO attainment without the overhead that compromises the serving throughput. Existing model-serving frameworks that leverage optimization techniques such as linear programming have exponential or cubical complexity, which limits the scalability to larger workloads and longer queues. QLM instead introduces request groups to reduce the input space of the optimization solvers, thus lowering the computational overhead and enabling scalability.
Design Principle #2: Handling multiple LSOs with inter dependencies. To attain model-serving latency SLOs, it is critical to translate the latency SLO to the appropriate backend LSO actions. QLM models these complex inter-relationships with a two-step approach. First, virtual queues enable the necessary abstraction to enable actions for multiple backend LSOs. Second, the global scheduler models the impact of ordering on multiple LSOs with a linear programming solver. We specifically prefer a linear programming solver over other optimization methods because it systematically considers various constraints introduced by multiple LSOs, SLO constraints, and waiting time estimates from the RWT estimator.
Design Principle #3: Handling heterogeneous models and hardware device configurations. LLM serving workloads consist of diverse model types with vastly different computational requirements, SLOs, and token length distributions. Hardware device configurations are also heterogeneous in terms of computing power, GPU memory capacity, and GPU-CPU memory bandwidth. To efficiently map LLM requests to the appropriate hardware resources, QLM’s global scheduler has to consider each device’s computing power, memory capacity, and memory bandwidth. The RWT estimator estimates this impact of heterogeneity for the global scheduler. The profiling costs for the RWT estimator are minimal, only a single batch run for a given combination of request group and GPU device is needed. Hence, QLM does not require significant training when adding new LLM models or GPU devices into the serving cluster.
4. Request Groups and Virtual Queues
In this section, we describe the concept of virtual queues and the process of classifying LLM requests into request groups and assigning request groups to virtual queues.
@@ 4. Request Groups and Virtual Queues
Request Group: Each request group is a collection of multiple requests that are relatively homogeneous, i.e., sharing similar performance demand or requirement characteristics. We identify that input/output token distributions, model type, and SLO values are sufficient for the RWT estimator (as explained in Section 6).
@@
By creating the abstraction of request groups and virtual queues, the ordering of request groups in a virtual queue allows QLM to configure actions for multiple downstream LSOs to attain latency SLOs for the LLM-serving requests in a scalable manner (described in Section 5).
@@
Request Group Creation. Request groups are created in two steps: (i) clustering similar requests based on definition 4.1, and (ii) splitting large request groups. Algorithm 1 describes the request group creation process. The parameters identified for the request grouping include model types, input/output token distribution, and SLO values. Grouped requests based on such parameters exhibit predictable request completion time distribution (compared to that of each individual request) as explained in Section 6. Additionally, we also limit the size of each request group to a small multiple ($\delta$) of batch size. We refer the reader to section 8.3 for the trade-off analysis between overhead and decision-making granularity: (1) Larger request groups would decrease the number of request groups and thus the overhead of the global scheduler; (2) However, restricting the size of request groups is beneficial as it allows for more fine-grained decisions. Since requests within a request group are relatively homogeneous, QLM treats the ordering of the requests within a group using a first-come-first-serve (FCFS) policy. Request groups are dequeued from the virtual queue when all requests complete execution.
@@
Handling New Incoming Requests. As new requests join the global queue, they are classified into the existing request groups, and the RWT estimator calculation is triggered to find out whether any SLOs are being violated. Upon any SLO violation, the global scheduler is called to reorder the request groups in the virtual queues to maximize SLO attainment given the current states (estimations).
@@
Fault Tolerance in Queue Management. QLM only stores a single replica of the requests and their metadata in the global queue, which avoids the need to maintain consistency between multiple queues. The global queue is implemented using a distributed message queue broker such as RabbitMQ (RabbitMQ, 2024) that provides the necessary replication, fault tolerance, and persistence mechanisms. The data structure that implements virtual queues records orderings of subsets of requests in the global queue. These virtual queues are implemented as lightweight data structures that maintain pointers or references to the actual requests stored in the global queue. By using virtual queues, QLM can achieve the following: (1) Fault Isolation: If an LLM serving instance fails, only the corresponding virtual queue is affected, and the remaining virtual queues can continue processing requests without interruption. Request groups from the lost virtual queue are assigned to other virtual queues using the global scheduler. (2) Consistency: Since the actual requests are stored in the global queue, virtual queues can be reconstructed or reassigned without compromising the consistency of the request data.
5. LLM Serving Operations
@@
The LLM serving instances serve requests from the corresponding virtual queue and execute backend LSO actions when necessary. The LSOs by themselves are merely action actuators, and the intelligence required to configure when and which action to set comes from the virtual queue ordering set by the global scheduler (as described in section 7).
@@
fig. 7 shows the four basic LSOs that QLM currently supports. A QLM agent resident on each LLM serving instance monitors the virtual queue ordering and converts it into LSO actions. When the virtual queue state changes or new requests are added, QLM agents initiate request pulling and load balancing. Similarly, when the head request group changes, the QLM agent initiates request eviction. Model swapping is initiated by the QLM agent for models at the head of the virtual queue. Each of these LSOs modifies the internal state of the LLM serving instance, which includes the running batch of requests, KV cache store, and model weights. Below, we describe each of these LSOs in detail and their action setup based on the virtual queue ordering.
@@
Figure 7. Basic LLM serving operations (LSOs) for an LLM-serving instance that a QLM agent manages.
[…]
Basic LLM serving operations (LSOs) for an LLM-serving instance that a QLM agent manages.
@@
Request Eviction ( 2 in fig. 7). As request pulling by itself may not be sufficient to enable the immediate serving of requests with low SLO values due to head-of-the-line blocking, QLM also supports request eviction. Request eviction is invoked when the RWT estimator detects an SLO violation, and the global scheduler replaces an existing request group by placing a request group at the head of the virtual queue. In request eviction, requests of the head request group are pulled into the running batch based on available capacity, and previously running requests are evicted (back) into the global queue. To prevent re-computation of KV cache of the lost request upon eviction, we migrate it to CPU memory instead. However, the GPU-to-CPU memory bandwidth is typically at least 10$\times$ less than the GPU memory bandwidth, and if the evicted request has a large KV cache, it leads to significant transfer overhead and consequent performance degradation. QLM hides this performance degradation using the asynchronous GPU memory copy available in most GPU programming libraries.
@@
Load Balancing ( 3 in fig. 7). As each LLM serving instance is associated with a separate virtual queue, the global scheduler’s assignment of request groups to a virtual queue inherently performs load balancing. Each instance would only pull from its associated virtual queue, thus ensuring the distribution of requests across all the serving instances. Note that QLM does not implement preemptive load balancing, i.e., once request groups start executing on an LLM serving instance, they cannot be migrated to another instance.
@@
Model Swapping ( 4 in fig. 7). Each LLM serving instance can serve multiple models by switching the underlying model weights and flushing out the KV cache. QLM assumes a two-tier hierarchy of memory and disk storage. Therefore, any model that needs to be served from the LLM model registry (located in the storage) has to undergo two distinct swaps: (1) Storage-CPU swapping:The model is first swapped from the LLM model registry to CPU memory, and (2) CPU-GPU swapping:The model located in the CPU memory is swapped into GPU memory for inferencing. QLM is able to decide the location (i.e. GPU mem, CPU mem or storage) of each model by checking the virtual queue order. The model for the head request of the virtual queue is currently active and should be placed in the GPU memory. Models present later in the virtual queue are warm and placed in the CPU memory until all the CPU memory is exhausted. The remaining models (cold models) are not swapped out from the LLM model registry (located in the storage).
@@
LSO Implementation. We implement the abovementioned LSOs on top of vLLM, a state-of-the-art LLM serving system. QLM agents are responsible for triggering each LSO action. Request pulling and load balancing are implemented by async pull calls to the virtual queues when there is spare token capacity on the LLM serving instance. Request eviction and GPU-CPU state swapping require instrumentation to the vLLM scheduler. In each iteration, the vLLM scheduler attempts to generate a new token for all the running requests and preempts any request that exceeds the total GPU capacity. At the end of an iteration, QLM agent checks if either request eviction or state swapping is required and performs the operation. For both operations, current requests are removed from the running batch to make space for the incoming requests at the head of the virtual queue. For request eviction, the scheduler uses an asynchronous GPU transfer of the KV cache. Model swapping into the GPU is implemented by changing the underlying model of the vLLM instance and flushing out the KV cache.
6. Request Waiting Time (RWT) Estimator
Table 1. Glossary of symbols used in the RWT estimator.
@@ 6. Request Waiting Time (RWT) Estimator
QLM leverages waiting time estimates from the RWT estimator for better utilization and decrease resource costs by making closer to optimal request queue reordering decisions. The RWT estimator uses a statistical approach to generate estimates of waiting and completion times for requests. To do so, the estimator first generates an upper bound on request completion time i.e. the actual request completion time would be lesser than the estimate. These completion times are then aggregated to generate completion time of the entire request group. Overall, the estimator is conservative when queue size is small (i.e. the estimate is higher than the actual waiting times) and the accuracy increases with increase in queue size.
The estimation process is explained in further detail below with variable definitions listed in table 1.
As shown in Equation 1, the total request completion time equals the sum of the waiting time ($W_{q}$), prefill time ($P$), and total decode time across all the output tokens ($D_{q}$) for a request $q$.
(1) $C_{q}=W_{q}+P+D_{q}$
Estimating Prefill Time. The prefill time $P$ is typically constant per model type when the input tokens are small as it is a highly parallel GPU-accelerated operation whose time increases minimally as the number of input tokens increases  [^2] Abnormally long context prompts beyond 10 standard deviations of production prompt distribution are not in the scope of QLM and we leave it for future research work.. Experiments show that the latency increase from additional input tokens is 100$\times$ less compared to the latency increase from each additional output token (Anyscale, 2024). Therefore, the major distribution terms that still remain to be estimated are the waiting time ($W_{q}$) and decode time ($D_{q}$).
Estimating Waiting Time. We consider the token generation throughput ($\Theta$) to be constant throughout the token generation process due to statistical averaging effects described in Appendix A.1. Therefore, the total waiting time for a single request can be represented by Equation 2 by dividing the number of tokens ahead ($\sum_{i=1}^{q-1}O_{i}$) in the queue by the token generation throughput ($\Theta$) where $i$ denotes each of the $q-1$ requests in the queue ahead of the request we model.
(2) $W_{q}=\sum_{i=1}^{q-1}\frac{O_{i}}{\Theta}$
Note that we do not know the number of output tokens ahead of time (that requires the knowledge of the output sequence for all requests in the waiting queue), so we model them as a distribution with the mean $\mu_{o}$ and standard deviation $\sigma_{o}$ fitted from the request input-output history dataset for the request group that the request $q$ belongs to. As $q$ becomes larger, the Central Limit Theorem (CLT) applies and the assumption of Normal distribution is accurate for any underlying request output token distribution. We further explain this in Appendix A.1.
(3) $\sum_{i=1}^{q-1}O_{i}\sim N((q-1)\mu_{o},(q-1)\sigma_{o}^{2})$
@@
(4) $D_{q}=O_{q}\times\epsilon\times d$
As we do not know the exact number of output tokens ($O_{q}$) in Equation 4 and the Normal distribution assumption from CLT does not apply for a single request, we approximate it using the maximum possible number of output tokens for the model. As the request queue grows, the waiting time $W_{q}$ dominates the completion time $C_{q}$ and the error introduced by the above heuristic reduces as shown in Figure 20. However, for short queues, we maintain the conservative estimate of decode time as it dominates the completion time.
If GPU memory was not a constraint, the decode steps would not be interrupted, and the total decode time would simply be the product of the number of output tokens ($O_{q}$) and decode time per output token ($d$). However, LLM serving systems cannot ensure this ideal behavior due to continuous batching. As requests are added continuously to the GPU’s running batch, some requests inevitably exceed the total GPU memory capacity limit and have to be temporarily preempted. This leads to inefficiency in the generation process that we capture with the inefficiency factor $\epsilon$, i.e., a constant multiplied by the decode time per token that captures the inefficiency associated with the generation process.
@@
(5) $C=\max_{q}C_{q}$
Offline Profiling. There are two independent profiling steps required for the RWT estimator: (a) Workload Profiling:samples multiple requests from the workload to generate a distribution for input and output tokens, and (b) Hardware Profiling:requires running the model with a single batch of requests on the specific GPU. Fixed variables associated with the model and hardware setup, such as the prefill time ($P$), inefficiency factor ($\epsilon$) and decode time per iteration ($d$) are obtained by directly logging these metrics from the LLM serving instance. In our implementation, we add these logging metrics directly into vLLM code.
7. Global Scheduler
The global scheduler is invoked by the RWT estimator when an SLO violation is likely to occur. Upon invocation, the global scheduler runs a linear programming model to reorder the virtual queues that decide underlying LSO actions to maximize SLO attainment. The global scheduler uses a linear program solver because it: (a) allows handling non-determinism by representing request group completion times as distributions, and (b) offers a systematic way to model various constraints associated with SLOs, model swapping times, and hardware heterogeneity. In this section, we present the linear programming model with its defined variables listed in table 2.
Table 2. Glossary of symbols used in the linear programming solver.